Broncos:

Which is a valid proportion? Three-fourths = StartFraction 21 over 28 EndFraction StartFraction 5 over 6 EndFraction = StartFraction 29 over 30 EndFraction StartFraction 4 over 7 EndFraction = StartFraction 16 over 49 EndFraction Two-fifths = StartFraction 12 over 15 EndFraction

Broncos:

Isabella worked 50 hours over the past two weeks, and she gets paid by the hour. During the first week of the two-week span, she worked 30 hours and got paid $285.00. How much did she get paid during the second week of the two-week span? $114.00 $190.00 $427.50 $475.50

Mercury:

old question but will respond so this can be closed starting w/ the first one 3/4 = 12/28 5/6 = 29/30 4/7 = 16/49 2/5 = 12/15 for each answer choice, simply test whether the left and right sides are equal. fast way would be to just divide numerator/denominator and see if both sides are still the same decimal

Mercury:

sescond one: Isabella worked 50 hours over the past two weeks, and she gets paid by the hour. During the first week of the two-week span, she worked 30 hours and got paid $285.00. How much did she get paid during the second week of the two-week span? $114.00 $190.00 $427.50 $475.50 the total amount worked is 50 hours. she worked 30 hours the first week. so she must have worked 50 - 30 = 20 hours the second week since the hourly rate is the same for both weeks (pay during second week)/(# of hours worked second week) = (pay during first week)/(# of hours worked first week) solve for second week's pay